Presentation is loading. Please wait.

Presentation is loading. Please wait.

FROST G9b FROST status update May 24 - June 13, 2010 1 Run Coordinators: Mike Dugger, Eugene Pasyuk, Barry Ritchie.

Similar presentations


Presentation on theme: "FROST G9b FROST status update May 24 - June 13, 2010 1 Run Coordinators: Mike Dugger, Eugene Pasyuk, Barry Ritchie."— Presentation transcript:

1 FROST G9b FROST status update May 24 - June 13, Run Coordinators: Mike Dugger, Eugene Pasyuk, Barry Ritchie

2 FROST Statistics and settings Circularly polarized beam Energy (GeV)Events (Billions) 35 Linearly polarized beam Edge (GeV)Events (Billions) (running now)Amorphous 0.6 Billion 2

3 FROST Summary (5/24 to 6/12) Shifts total: 63 Shifts with no major problems (>6 h production): 38 Re-polarization downtime (takes ~1.5 shifts, but coupled with beam studies, so x 0.5): ~ 1.5 Lost beamtime tied to specific principal sources: –CHL cold box trips: ~3 –DAQ crashes, issues: ~3 (integrated) –DC HV, fuses: ~2 (integrated) –MCC. “hurricane”: ~1.5 Remainder: Other miscellaneous machine trips, beam interruptions, goniometer changes and issues, etc. (~14) 3

4 FROST Monday May 24 OWLPRODUCTION DATA OWL: PRODUCTION DATA DAYPRODUCTION DATA DAY: PRODUCTION DATA (except 1.5 hours of no beam) SWINGPRODUCTION DATA SWING: PRODUCTION DATA (except ~1 hour of problems) 23:00 → DC HV crash required entry into the hall. 4

5 FROST Tuesday May 25 OWL OWL: DC HV problems (except ~ 2 hrs. of production data) 00:00 → Rebooted CAEN crate 1 01:40 → Reset VME DC controller 03:13 → V0 values noticed to be set to zero 05:40 → Replaced board in crate 4, slot 2 PRODUCTION DATA 06:10 → Start of PRODUCTION DATA DAY DAY: Beam studies and target polarization SWING SWING: 16:00-19:30 → Finished polarization (92.5 %) 19:50-23:50 → Beam has unacceptable current and position fluctuations PRODUCTION DATA 23:50 → Start of PRODUCTION DATA 5

6 FROST Wednesday May 26 OWLPRODUCTION DATA OWL: PRODUCTION DATA (except ~ 1hr of DC and beam ) 00:44-01:09 → DC R1S4B HV problem 04:34-05:15 → No beam DAY DAY: DC HV work in the hall. Must use EPICS to change HV values, or else all values will need to be put in by hand SWING SWING: 17:15 → Beam sent to hall 17:55-19:30 → Goniometer GUI crash. Trouble connecting device to GUI. DAQ crash that required expert help. 19:30-23:35 → Data of questionable quality 22:00 → L2 trigger problem discovered. DAQ expert consulted 23:35 → DAQ expert requests controlled access to the hall 6

7 FROST Thursday May 27 OWLPRODUCTION DATA OWL: PRODUCTION DATA (except ~ 1.5 hours of DAQ work) NOTE: Dead time is now about 50% of typical. DAYPRODUCTION DATA DAY: PRODUCTION DATA (except ~1.8 hr of no beam) 14:18-16:00 → No beam (2k cold box trip). SWING SWING: 16:00-18:50 → No beam PRODUCTION DATA 19:18-24:00 → PRODUCTION DATA 7

8 FROST Friday May 28 OWLPRODUCTION DATA OWL: PRODUCTION DATA DAY DAY: PRODUCTION DATA (except ~3.5 hours) 09:08-11:20 → Computer center attacked the ROCs. They forgot to mask the Hall-B ip addresses when doing a security check. This caused us to have a controlled access to flash some ROCs 13:45-14:30 → Controlled access for DC experts change trip point for R2S5 low voltage. 14:53-15:20 → Controlled access: DC expert investigates tripped HV channel (R3S2AXG0916). Determined that the problem was a guard wire that can be ignored for now. SWINGPRODUCTION DATA SWING: PRODUCTION DATA 8

9 FROST Saturday, May 29 OWLPRODUCTION DATA OWL: PRODUCTION DATA DAYPRODUCTION DATA DAY: PRODUCTION DATA (except ~2 hr of DAQ and 20 minutes of no beam) 16:00 → Eugene takes over as run coordinator. Swing: Smooth running 9

10 FROST Sunday, May 30 Smooth running all day long! 10

11 FROST Monday, May 31 Owl: production running. Blown fuse R2S1. Day: Smooth running Swing: Smooth running 11

12 FROST Tuesday, June 1 Owl: Smooth running Day: Target polarization. R2S1 fixed Swing: resume data taking around 20:00 12

13 FROST Wednesday, June 2 Smooth running all day long! 13

14 FROST Thursday, June 3 Owl: smooth running Day: Hurricane at MCC Swing: recovery from the hurricane. Beam restored by midnight 14

15 FROST Friday, June 4 Owl; Taking data. Blown fuse R2S6 Day: Taking data. 09:30-10:30 access to replace the fuse. 14:00-15:30 Test of active collimator as a device for orbit lock. Needs more studies. Swing: 16:46 cold box trip no beam for the rest of the shift 15

16 FROST Saturday, June 5 Owl: beam is back at 01:00. Start production. At 5:30 goniometer moves on its own! Restart GUI. Day: production running. 13:30 cold box tripped again Swing: no beam 16

17 FROST Sunday, June 6 Owl: Beam is back at 06:30. resume data taking Day: taking data. DAQ problems (1.5 hours) Swing: production data 17

18 FROST Monday, June 7 Owl shift: Production running Day shift: BGR takes over RC, smooth running Swing shift: Smooth running 18

19 FROST Tagger calibration 19

20 FROST Start counter 20

21 FROST TOF 21

22 FROST First look at target asymmetry T Use data from 2 “uncalibrated” runs (62298, 62375) 650 < E  < 1200 MeV Assume  + p →   + X 0.9 < cos  cm < 1.0 (meson angle) 0.9 < Mass X < 1.0 GeV 22 Analysis by Mike Dugger

23 FROST First look at target asymmetry T 23 Mike Dugger

24 FROST First look at target asymmetry T 24 Fit target offset angle No background subtraction Target offset = 60 degrees Background subtraction 650 < E  < 1200 MeV, 0.9 < cos  cm < 1.0 (meson angle), 0.9 < Mass X (final state  + X)< 1.0 GeV Mike Dugger

25 FROST Summary The g9b run is going very well so far –FROST target is performing superbly, reliably –CLAS and tagger are working steadily –Accelerator is operating well –Shift workers are doing a great job –Initial calibrations progressing nicely –First peek at data portends lots of physics results ahead 25


Download ppt "FROST G9b FROST status update May 24 - June 13, 2010 1 Run Coordinators: Mike Dugger, Eugene Pasyuk, Barry Ritchie."

Similar presentations


Ads by Google